# A lower bound on the number of edges in DP-critical graphs

### v1

@@ 3 Properties of GDP-forests
$\widetilde{B}$ is $t$-regular for some $t\geq 2$, and $j=qt\geq 2q$. (12)
For each $v\in(V(B)-u)\cap V_{k-1}(T)$, $v$ contributes $\alpha(k-1-j)+0-1$ to $\Phi_{k}(v)$, and for each $v\in(V(B)-u)\cap V_{k-1}^{-}(T)$, $v$ contributes at least $1$ to $\Phi_{k}(v)$. For two adjacent vertices $x,y\in V(B)$, the $q$ edges in $E_{B}(x,y)$ contribute $q-1$ to $\Phi_{k}(T)-\Phi_{k}(T^{\prime})$ due to the increase in $m(T)$. Moreover, for each neighbor $x\in V(B)$ of $u$, the $q$ edges in $E_{G}(u,x)$ contribute an additional $-\alpha q$ to $\Phi_{k}(T)-\Phi_{k}(T^{\prime})$ due to the decrease in $h(u)-d_{T}(u)$. Note that $s\geq\frac{j}{q}$, and $B$ has at least $(s+1)q$ edges, so $m(B)\geq(s+1)(q-1)$.
@@
Otherwise, $j=k-3$, so each vertex $v\in V(B)-u$ contributes at least $\min\{\alpha(k-1-j)-1,1\}=\min\{2\alpha-1,1\}=2\alpha-1$ to $\Phi_{k}(T)$. If $q\geq 2$, then $j=qt\geq 4$. Then, by (12), $k=3+j\geq 7.$ If $s=2$, then $\widetilde{B}$ is a $3$-cycle, so $j$ is even and $m(B)=\frac{3j}{2}-3\geq j-1$ since $j\geq 4$. If $s\geq 3$, then $m(B)=\frac{j-t}{2}(s+1)\geq\frac{j}{4}(3+1)=j$. Hence, $m(B)\geq j-1$.
@@
The remaining possibility is that $j=k-3$ and $q=1$, which means that $B=K_{k-2}$. Then $d_{T^{\prime}}(u)=d_{T}(u)-j\leq 2$, so write $N_{T^{\prime}}(u)=\{u_{1},u_{2}\}$ (possibly, $u_{2}=u_{1}$).
@@
    each $v\in(V(B)-u)\cap V_{k-1}(T)$ contributes $12\alpha-$ 1 to $\alpha(h(v)-d_{T}(v))-|V_{k-1}(T)\cap\{v\}|$,
so the vertices in $V(B)-u$ collectively contribute at least $(k-3)\min\{1,2\alpha-1\}=(k-3)(2\alpha-1)$ to $\Phi_{k}(T)$.
@@
Case 1: Assume $u_{1}=u_{2}$, so $N_{T^{\prime}}(u)=\{u_{1}\}$. Note that $u$ contributes no less than $-1$ to $\Phi_{k}(T)$.
@@
Case 2.1: $|V(B_{0})|\geq 4$, so $B_{0}$ is a cycle of length at least $4$, and $u_{1}u_{2}$ is not an edge. Obtain $T_{1}$ from $T^{\prime\prime}$ by adding the edge $u_{1}u_{2}$, so $T_{1}$ is a good tree. Then $u$ contributes $-1$ to $\Phi_{k}(T)-\Phi_{k}(T_{1})$. Together with (*),
@@
Case 2.2: $|V(B_{0})|=3$ and $d_{T}(u_{1})=2$. Obtain $T_{2}$ from $T^{\prime\prime}$ by deleting $u_{1}$, so that $T_{2}$ is a good tree. Then $d_{T}(u_{2})-d_{T_{2}}(u_{2})=2$, and $u_{1}$ contributes to $\Phi_{k}(T)-\Phi_{k}(T_{2})$ either $\alpha(k-1-2)-1$ (when $h(u_{1})=k-1$) or at least $1$ (when $h(u_{1})\leq k-2$). As in Case 2.1, $u$ contributes $-1$ to $\Phi_{k}(T)-\Phi_{k}(T_{2})$. Together with (*),
@@
Case 2.3: $|V(B_{0})|=3$ and both $u_{1}$ and $u_{2}$ are cut vertices. For $i\in\{1,2\}$, let $B_{i}$ be a block of $T$ containing $u_{i}$ and distinct from $B_{0}$. Since $B$ is peripheral, at least one of $B_{1},B_{2}$ also is peripheral, say $B_{1}$. Then by the choice of $B$, $B_{1}$ is $(k-3)$-regular, and so by the previous cases, $B_{1}=K_{k-2}$. In this case, $T_{3}=T-V(B)-V(B_{1})$ is a good tree. Then $d_{T}(u_{2})-d_{T_{3}}(u_{2})=2$, and each of $u,u_{1}$ contributes $-1$ to $\Phi_{k}(T)-\Phi_{k}(T_{2})$. Together with (*) for both $B$ and $B_{1}$,
@@
4 Properties of a minimum counterexample $G$
For the rest of the paper, we fix a counterexample $G$, which is a loopless multigraph with a function $h:V(G)\to\{0,\ldots,k-1\}$, to Theorem 2.1 with minimum $|V(G)|$. We also fix an $h$-cover $(H,L)$ of $G$ such that $G$ does not have an $(H,L)$-coloring. For every $uv\in E(G)$, we assume $E_{H}(L(u),L(v))$ is the union of $|E_{G}(u,v)|$ maximal matchings. We write $L(v)=\{1_{v},\dots,t_{v}\}$, where $t=|L(v)|$.
4.1 General observations
@@
For each induced subgraph $X\subseteq G$, $X$ is not DP $(h-d_{G}+d_{X})$-colorable.
@@
Let $(H,L)$ be an $h$-cover of $G$ for which $G$ admits no $(H,L)$-coloring. As $G$ is $h$-minimal, $G-V(X)$ has an $(H,L)$-coloring $f$. Now, let $H^{\prime}$ be the restriction of $H$ to $L(V(X))$, and for each $x\in V(X)$, let $L^{\prime}(x)\subseteq L(x)$ consist of the colors in $L(x)$ that are not adjacent to a color of $f$. As $G$ has no $(H,L)$-coloring, $f$ cannot be extended to $X$; therefore, $X$ has no $(H^{\prime},L^{\prime})$-coloring. As $|L^{\prime}(x)|\geq h(x)-d_{G}(x)+d_{X}(x)$ for each $x\in V(X)$, $X$ is not DP $(h-d_{G}+d_{X})$-colorable. ∎
@@
If $U\subseteq V(G)$ is a set of low vertices, then each block of $G[U]$ is isomorphic to $K_{t}^{s}$ or $C_{t}^{s}$, for some $t$ and $s$.
@@
Suppose that $G[U]$ contains a block $B$ that is isomorphic to neither $K_{t}^{s}$ nor $C_{t}^{s}$. Let $(H,L)$ be an $h$-cover of $G$ for which $G$ has no $(H,L)$-coloring. As $G$ is $h$-minimal, $G-V(B)$ has an $(H,L)$-coloring $f$. Now, for each $v\in V(B)$, let $L^{\prime}(v)\subseteq L(v)$ consist of the colors in $L(v)$ with no neighbor in $f$. Each $v\in V(B)$ is low, so $|L^{\prime}(v)|\geq d_{B}(v)$. Then, $B$ has an $(H,L^{\prime})$-coloring by Theorem 1.7, and thus $G$ is $(H,L)$-colorable, a contradiction. ∎
@@
Let $G^{\prime}=G-S$. For each vertex $v\in V(G^{\prime})$, let $h^{\prime}(v)=\max\{0,h(v)-d_{G}(v)+d_{G^{\prime}}(v)\}=\max\{0,h(v)-|E_{G}(v,S)|\}$. By Observation 4.2, $G^{\prime}$ is not DP $h^{\prime}$-colorable. Therefore, there exists $U\subseteq V(G^{\prime})$ for which $G^{\prime}[U]$ has a spanning $h^{\prime}$-minimal subgraph. As $G$ is $h$-minimal, it follows that $(G,h)$ and $(G^{\prime},h^{\prime})$ do not agree on $U$; therefore, $U$ contains a neighbor $u$ of $S$. As $h^{\prime}(u)<k-1$, $\rho_{G^{\prime},h^{\prime}}(U)\leq-2$.
@@
$\rho_{G,h}(S^{\prime})\leq\rho_{G,h}(U)+\rho_{G,h}(S)-2\lambda\ell\leq-2+\ell(\lambda+2)-\ell(2\lambda)+\rho_{G,h}(S)=-2-\ell(\lambda-2)+\rho_{G,h}(S).$ (13)
@@
$\rho_{G,h}(S^{\prime})\leq-2-\ell(\lambda-2)+j(\lambda-2)=-2+(j-\ell)(\lambda-2).$ (14)
@@
Now, since $|E_{G}(S,U)|=\ell$ and $|E_{G}(S,\overline{S})|\geq j$, it follows that $|E_{G}(S^{\prime},\overline{S^{\prime}})|\geq|E_{G}(S,V(G^{\prime})\setminus U)|\geq j-\ell$. As $11\leq j-\ell\leq j-$ 1, the minimality of $j$ tells us that $\rho_{G,h}(S^{\prime})>(j-\ell)(\lambda-2)$, contradicting (14). ∎
@@
If $11\leq|U|\leq|V(G)|-$ 1, then $\rho_{h}(U)\geq\lambda-1$ by Lemma 4.5. If $|U|=|V(G)|$, then $\rho_{h}(U)=\rho_{h}(G)\geq-1$ by our assumption that $G$ is a counterexample to Theorem 2.1. ∎
@@
Suppose that $E_{G}(S,\overline{S})=\{x_{0}y_{0}\}$, where $x_{0}\in S$ and $y_{0}\in\overline{S}$. If we define $h^{\prime}:V(S)\to\mathbb{N}\cup\{0\}$ so that $h^{\prime}(x_{0})=h(x_{0})-1$ and $h^{\prime}(v)=h(v)$ for each $v\in S\setminus\{x_{0}\}$, Observation 4.2 implies that $G[S]$ is not $h^{\prime}$-colorable, and hence $S$ has a subset $X$ for which $G[X]$ has a spanning $h^{\prime}$-minimal subgraph. As $G$ is $h$-minimal, $h$ and $h^{\prime}$ do not agree on $X$, so $x_{0}\in X$. As $h^{\prime}(x_{0})<k-1$, $\rho_{G,h^{\prime}}(X)\leq-2$. As $\rho_{h}(x_{0})-\rho_{h^{\prime}}(x_{0})\leq\lambda+2$, we have $\rho_{G,h}(X)\leq\rho_{G,h^{\prime}}(X)+(\lambda+2)\leq-2+(\lambda+2)=\lambda$. By a symmetric argument, $\overline{S}$ has a vertex subset $Y$ containing $y_{0}$ for which $\rho_{G,h}(Y)\leq\lambda$.
@@
If $h(v)\in\{2,\ldots,\frac{\lambda-1}{2}\}$, then $v$ is low.
@@
We have $\rho_{h}(v)=h(v)\lambda-1\leq\lceil\frac{h(v)\lambda-1}{\lambda-2}\rceil(\lambda-2)$. Then, Lemma 4.5 implies that $d(v)\leq\lceil\frac{h(v)\lambda-1}{\lambda-2}\rceil-1=h(v)+\lceil\frac{2h(v)-1}{\lambda-2}\rceil-1=h(v)$. ∎
@@
Now, let $G^{\prime}=G-v$. Choose a color $c\in L(v)$, and obtain $H^{\prime}$ from $H$ by deleting $L(v)\cup N_{H}(c)$. We may assume without loss of generality that $|L(w)\cap N_{H}(c)|\geq 2$. Let $L^{\prime}(w)=L(w)\cap V(H^{\prime})$, and let $L^{\prime}(x)=L(x)$ for each $x\in V(G^{\prime})-w$. If $G^{\prime}$ has an $(H^{\prime},L^{\prime})$-coloring $f^{\prime}$, then we can extend it to $G$ by letting $f^{\prime}(v)=c$. So assume $G^{\prime}$ has no $(H^{\prime},L^{\prime})$-coloring. Then there exists a subset $U^{\prime}\subseteq V(G^{\prime})$ for which $G^{\prime}[U^{\prime}]$ has a spanning $h^{\prime}$-minimal subgraph. As $G$ is $h$-minimal and $U^{\prime}\subsetneq V(G)$, it follows that $h^{\prime}$ and $h$ do not agree on some vertex of $U^{\prime}$; therefore, $w\in U^{\prime}$. As $h^{\prime}(w)<k-1$, it follows from the minimality of $G$ that $\rho_{G^{\prime},h^{\prime}}(U^{\prime})\leq-2$, so that $\rho_{G,h}(U^{\prime})\leq-2+\lambda+(\lambda+2)=2\lambda$. Then,
@@
Obtain $G^{\prime}$ from $G-v$ by adding an edge $uw$. Also, let $(H^{\prime},L)$ be an $h$-cover of $G^{\prime}$ obtained by deleting $L(v)$ and then adding the matching $\{1_{u}2_{w},2_{u}1_{w}\}$. (Here, we consider $h$ restricted to $V(G^{\prime})$.) If $G^{\prime}$ has an $(H^{\prime},L)$-coloring $f$, then we can extend $f$ to $v$, because if $f(u)$ forbids the color $i_{v}$ for $v$, then (due to our new edge $uw$) $f(w)$ does not forbid the color $(3-i)_{v}$ for $v$. Thus, $G^{\prime}$ has an $h$-minimal subgraph $G^{\prime\prime}$ containing $u$ and $w$. Let $U^{\prime}=V(G^{\prime\prime})$.
As $|V(G^{\prime})|<|V(G)|$, it follows from our induction hypothesis that $\rho_{G^{\prime},h}(U^{\prime})\leq-2$ or $G^{\prime\prime}$ is a copy of $K_{k}$. If $\rho_{G^{\prime},h}(U^{\prime})\leq-2$, then $G[U^{\prime}\cup\{v\}]$ is constructed by deleting an edge $uw$ from $G^{\prime}$, adding a vertex $v$ with $h(v)=2$, and adding two edges $vu$ and $vw$; hence,
$\rho_{G,h}(U^{\prime}\cup\{v\})\leq-2+(2\lambda-1)-2(2\lambda)+(2\lambda+1)=-2,$
contradicting Observation 4.6. Therefore, $G^{\prime\prime}$ is a copy of $K_{k}$, and $h(x)=k-1$ for each $x\in U^{\prime}$.
Then, writing $U=U^{\prime}\cup\{v\}$,
$\rho_{G,h}(U)=k+(2\lambda-1)-2(2\lambda)+2\lambda=k-1.$

### v2

@@ 3. Properties of GDP-forests
(13) $\widetilde{B}$ is $t$-regular for some $t\geq 2$, and $j=qt\geq 2q$.
For each $\in(V(B)-u)\cap V_{k-1}(T), vv$ contributes $\alpha(k-1-j)+0-1$ to $\Phi_{k}(v)$, and for each $\in(V(B)-u)\cap V_{k-1}^{-}(T), vv$ contributes at least $1$ to $\Phi_{k}(v)$. For two adjacent vertices $x,y\in V(B)$, the $q$ edges in $E_{B}(x,y)$ contribute q1q-1 to $\Phi_{k}(T)-\Phi_{k}(T^{\prime})$ due to the increase in $m(T)$. Moreover, for each neighbor $x\in V(B)$ of $u$, the $q$ edges in $E_{G}(u,x)$ contribute an additional $\alpha q$ to $\Phi_{k}(T)-\Phi_{k}(T^{\prime})$ due to the decrease in $h(u)-d_{T}(u)$. Note that $s\geq\frac{j}{q}$, and $B$ has at least $(s+1)q$ edges, so $m(B)\geq(s+1)(q-1)$.
@@
Otherwise, j=k3j=k-3, so each vertex $v\in V(B)-u$ contributes at least $\min\{\alpha(k-1-j)-1,1\}=\min\{2\alpha-1,1\}=2\alpha-1$ to $\Phi_{k}(T)$. If $q\geq 2$, then $j=qt\geq 4$. Then, by (13), $k=3+j\geq 7.$ If $s=2$, then $\widetilde{B}$ is a $3$-cycle, so $j$ is even and $m(B)=\frac{3j}{2}-3\geq j-1$ since $j\geq 4$. If $s\geq 3$, then $m(B)=\frac{j-t}{2}(s+1)\geq\frac{j}{4}(3+1)=j$. Hence, $m(B)\geq j-1$.
@@
The remaining possibility is that j=k3j=k-3 and $q=1$, which means that $B=K_{k-2}$ or $k=5$ and $B$ is a cycle. In the latter case, each $v\in T^{\prime}$ contributes at least $1\cdot(3-2)=1$, since $\alpha=1,h(v)\geq 3$, and $j=2$. Thus $\Phi_{k}(T)-\Phi_{k}(T^{\prime})>0$. In the former case, we have $d_{T^{\prime}}(u)=d_{T}(u)-j\leq 2$, so write $N_{T^{\prime}}(u)=\{u_{1},u_{2}\}$ (possibly, $u_{2}=u_{1}$).
@@
    each $v\in(V(B)-u)\cap V_{k-1}(T)$ contributes $2\alpha-1$ to $\alpha(h(v)-d_{T}(v))-|V_{k-1}(T)\cap\{v\}|$,
so the vertices in V(B)uV(B)-u collectively contribute at least $(k-3)\min\{1,2\alpha-1\}=(k-3)(2\alpha-1)$ to $\Phi_{k}(T)$.
@@
Case 1: Assume $u_{1}=u_{2}$, so $N_{T^{\prime}}(u)=\{u_{1}\}$. Note that $u$ contributes no less than 1-1 to $\Phi_{k}(T)$.
@@
Case 2.1: $|V(B_{0})|\geq 4$, so $B_{0}$ is a cycle of length at least $4$, and $u_{1}u_{2}$ is not an edge. Obtain $T_{1}$ from $T^{\prime\prime}$ by adding the edge $u_{1}u_{2}$, so $T_{1}$ is a good tree. Then $u$ contributes 1-1 to $\Phi_{k}(T)-\Phi_{k}(T_{1})$. Together with (*),
@@
Case 2.2: $|V(B_{0})|=3$ and $d_{T}(u_{1})=2$. Obtain $T_{2}$ from $T^{\prime\prime}$ by deleting $u_{1}$, so that $T_{2}$ is a good tree. Then $d_{T}(u_{2})-d_{T_{2}}(u_{2})=2$, and $u_{1}$ contributes to $\Phi_{k}(T)-\Phi_{k}(T_{2})$ either $\alpha(k-1-2)-1$ (when $h(u_{1})=k-1$) or at least $1$ (when $h(u_{1})\leq k-2$). As in Case 2.1, $u$ contributes 1-1 to $\Phi_{k}(T)-\Phi_{k}(T_{2})$. Together with (*),
@@
Case 2.3: $|V(B_{0})|=3$ and both $u_{1}$ and $u_{2}$ are cut vertices. For $i\in\{1,2\}$, let $B_{i}$ be a block of $T$ containing $u_{i}$ and distinct from $B_{0}$. Since $B$ is peripheral, at least one of $B_{1},B_{2}$ is also peripheral, say $B_{1}$. Then by the choice of $B$, $B_{1}$ is (k3)(k-3)-regular, and so by the previous cases, $B_{1}=K_{k-2}$. In this case, $T_{3}=T-V(B)-V(B_{1})$ is a good tree. Then $d_{T}(u_{2})-d_{T_{3}}(u_{2})=2$, and each of $u,u_{1}$ contributes 1-1 to $\Phi_{k}(T)-\Phi_{k}(T_{2})$. Together with (*) for both $B$ and $B_{1}$,
@@
4. Properties of a minimum counterexample $G$
Suppose that Theorem 2.1 does not hold for some $k\geq 5$. Then there are a multigraph $G$, a function $h:V(G)\to[0,k-1]$, and an $h$-cover $(H,L)$ such that
  1. (A)
    $G$ has no $(H,L)$-coloring, but every proper subgraph $G^{\prime}$ of $G$ is DP $h|_{V(G^{\prime})}$-colorable, and
  2. (B)
    none of the claims (1)–(3) of Theorem 2.1 holds for $G$ and $h$.
Among such quadruples $(G,h,H,L)$, choose one with minimum $|G|$ and fix it for the rest of the paper.
For every $uv\in E(G)$, we assume $E_{H}(L(u),L(v))$ is the union of $|E_{G}(u,v)|$ maximal matchings. We write $L(v)=\{1_{v},2_{v},\dots,t_{v}\}$, where $t=|L(v)|$.
4.1. General observations
@@
For each induced subgraph $\subseteq G, XX$ is not DP $(h-d_{G}+d_{X})$-colorable.
@@
As $G$ is $h$-minimal, GV(X)G-V(X) has an $(H,L)$-coloring $\varphi$. For each $x\in V(X)$, let $L^{\prime}(x)\subseteq L(x)$ consist of the colors of $L(x)$ that are not adjacent to a color of $\varphi$. Then, let $H^{\prime}$ be the subgraph of $H$ induced by the set $\bigcup_{x\in V(X)}L^{\prime}(x)$. As $G$ has no $(H,L)$-coloring, $\varphi$ cannot be extended to $X$; therefore, $X$ has no $(H^{\prime},L^{\prime})$-coloring. As $|L^{\prime}(x)|\geq h(x)-d_{G}(x)+d_{X}(x)$ for each $x\in V(X)$, $X$ is not DP $(h-d_{G}+d_{X})$-colorable. ∎
@@
If $U\subseteq V(G)$ is a set of low vertices, then each block of $G[U]$ is isomorphic to $sK_{t}$ or $sC_{t}$, for some $t$ and $s$.
@@
Suppose that $G[U]$ contains a block $B$ that is isomorphic to neither $sK_{t}$ nor $sC_{t}$. Let $(H,L)$ be an $h$-cover of $G$ for which $G$ has no $(H,L)$-coloring. As $G$ is $h$-minimal, GV(B)G-V(B) has an $(H,L)$-coloring $\varphi$. Now, for each $v\in V(B)$, let $L^{\prime}(v)\subseteq L(v)$ consist of the colors in $L(v)$ with no neighbor in $\varphi$. Each $v\in V(B)$ is low, so $|L^{\prime}(v)|\geq d_{B}(v)$. Then, $B$ has an $(H,L^{\prime})$-coloring by Theorem 1.5, and thus $G$ is $(H,L)$-colorable, a contradiction. ∎
@@
Let $G^{\prime}=G-S$. For each vertex $v\in V(G^{\prime})$, let $h^{\prime}(v)=\max\{0,h(v)-d_{G}(v)+d_{G^{\prime}}(v)\}=\max\{0,h(v)-|E_{G}(v,S)|\}$. By Observation 4.2, $G^{\prime}$ is not DP $h^{\prime}$-colorable. Therefore, there exists $U\subseteq V(G^{\prime})$ for which $G^{\prime}[U]$ has a spanning $h^{\prime}$-minimal subgraph. As $G$ is $h$-minimal, it follows that $(G,h)$ and $(G^{\prime},h^{\prime})$ do not agree on $U$; therefore, $U$ contains a neighbor $u$ of $S$. As $h^{\prime}(u)<k-1$, and as $G$ contains no $K_{k}$ subgraph, $\rho_{G^{\prime},h^{\prime}}(U)\leq-2$.
@@
(14) $\rho_{G,h}(S^{\prime})\leq\rho_{G,h}(U)+\rho_{G,h}(S)-2\lambda\ell\leq-2+\ell(\lambda+2)-\ell(2\lambda)+\rho_{G,h}(S)=-2-\ell(\lambda-2)+\rho_{G,h}(S).$
@@
(15) $\rho_{G,h}(S^{\prime})\leq-2-\ell(\lambda-2)+j(\lambda-2)=-2+(j-\ell)(\lambda-2).$
@@
Now, since $|E_{G}(S,U)|=\ell$ and $|E_{G}(S,\overline{S})|\geq j$, it follows that $|E_{G}(S^{\prime},\overline{S^{\prime}})|\geq|E_{G}(S,V(G^{\prime})\setminus U)|\geq j$-$\ell$. As $1\leq j-\ell\leq j-1$, the minimality of $j$ tells us that $\rho_{G,h}(S^{\prime})>(j-\ell)(\lambda-2)$, contradicting (15). ∎
@@
If $1\leq|U|\leq|V(G)|-1$, then $\rho_{h}(U)\geq\lambda-1$ by Lemma 4.5. If $|U|=|V(G)|$, then $\rho_{h}(U)=\rho_{h}(G)\geq-1$ by our assumption that $G$ is a counterexample to Theorem 2.1. ∎
@@
Suppose that $E_{G}(S,\overline{S})=\{x_{0}y_{0}\}$, where $x_{0}\in S$ and $y_{0}\in\overline{S}$. If we define $h^{\prime}:V(S)\to\mathbb{N}\cup\{0\}$ so that $h^{\prime}(x_{0})=h(x_{0})-1$ and $h^{\prime}(v)=h(v)$ for each $v\in S\setminus\{x_{0}\}$, Observation 4.2 implies that $G[S]$ is not $h^{\prime}$-colorable, and hence $S$ has a subset $X$ for which $G[X]$ has a spanning $h^{\prime}$-minimal subgraph. As $G$ is $h$-minimal, $h$ and $h^{\prime}$ do not agree on $X$, so $x_{0}\in X$. As $h^{\prime}(x_{0})<k-1$, and as $G$ contains no $K_{k}$ subgraph, $\rho_{G,h^{\prime}}(X)\leq-2$. As $\rho_{h}(x_{0})-\rho_{h^{\prime}}(x_{0})\leq\lambda+2$, we have $\rho_{G,h}(X)\leq\rho_{G,h^{\prime}}(X)+(\lambda+2)\leq-2+(\lambda+2)=\lambda$. By a symmetric argument, $\overline{S}$ has a vertex subset $Y$ containing $y_{0}$ for which $\rho_{G,h}(Y)\leq\lambda$.
@@
If $h(v)\in[2,\ldots,\frac{\lambda-1}{2}]$, then $v$ is low.
@@
We have $\rho_{h}(v)=h(v)\lambda-1\leq\lceil\frac{h(v)\lambda-1}{\lambda-2}\rceil(\lambda-2)$. Then, Lemma 4.5 implies that
$d(v)\leq\left\lceil\frac{h(v)\lambda-1}{\lambda-2}\right\rceil-1=h(v)+\left\lceil\frac{2h(v)-1}{\lambda-2}\right\rceil-1=h(v).$
@@
Now, let $G^{\prime}=G-v$. Choose a color $c\in L(v)$, and obtain $H^{\prime}$ from $H$ by deleting $L(v)\cup N_{H}(c)$. We may assume without loss of generality that $|L(w)\cap N_{H}(c)|\geq 2$. Let $L^{\prime}(w)=L(w)\cap V(H^{\prime})$, and let $L^{\prime}(x)=L(x)$ for each $x\in V(G^{\prime})-w$. If $G^{\prime}$ has an $(H^{\prime},L^{\prime})$-coloring $\varphi^{\prime}$, then we can extend it to $G$ by letting $\varphi^{\prime}(v)=c$. So assume $G^{\prime}$ has no $(H^{\prime},L^{\prime})$-coloring. Then there exists a subset $U^{\prime}\subseteq V(G^{\prime})$ for which $G^{\prime}[U^{\prime}]$ has a spanning $h^{\prime}$-minimal subgraph. As $G$ is $h$-minimal and $U^{\prime}\subsetneq V(G)$, it follows that $h^{\prime}$ and $h$ do not agree on some vertex of $U^{\prime}$; therefore, $w\in U^{\prime}$. As $h^{\prime}(w)\leq k-3$, it follows from the minimality of $G$ that $\rho_{G^{\prime},h^{\prime}}(U^{\prime})\leq-2$, so that $\rho_{G,h}(U^{\prime})\leq-2+\lambda+(\lambda+2)=2\lambda$. Then,
@@
Obtain $G^{\prime}$ from GvG-v by adding an edge $uw$. Also, let $(H^{\prime},L)$ be an $h$-cover of $G^{\prime}$ obtained by deleting $L(v)$ and then adding the matching $\{1_{u}2_{w},2_{u}1_{w}\}$. (Here, we consider $h$ restricted to $V(G^{\prime})$.) If $G^{\prime}$ has an $(H^{\prime},L)$-coloring $\varphi$, then we can extend $\varphi$ to $v$, because if $\varphi(u)$ forbids the color $i_{v}$ for $v$, then (due to our new edge $uw$) $\varphi(w)$ does not forbid the color $(3-i)_{v}$ for $v$. Thus, $G^{\prime}$ has an $h$-minimal subgraph $G^{\prime\prime}$ containing $u$ and $w$. Let $U=V(G^{\prime\prime})$.
As $|V(G^{\prime})|<|V(G)|$, it follows from our induction hypothesis that $\rho_{G^{\prime},h}(U)\leq-2$ or $G^{\prime\prime}$ is exceptional. We consider several cases.
First, suppose that $k=5$, $h(x)=4$ for each $x\in U$, and that $G^{\prime\prime}$ is a copy of $4K_{2}$ or $2C_{t}$ for some integer $t$. Since each adjacent pair of vertices in $G^{\prime\prime}$ has edge multiplicity at least $2$, and as $u,w\in U$, we have $uw\in E(G)$. Furthermore, as $G^{\prime\prime}$ is exceptional, $\rho_{G^{\prime\prime},h}(U)\leq 0$. Therefore,
$\rho_{G,h}(U\cup\{v\})\leq 0+(2\lambda+1)-2(2\lambda)+(2\lambda-1)=0.$
Thus, Observation 4.7 implies that $V(G)=U\cup\{v\}$. Furthermore, Observation 4.6 implies that $h(x)=4$ for each $x\in U$. Then, $G$ is DP $h$-colorable by Theorem 1.5, a contradiction.
Next, suppose that $\rho_{G^{\prime},h}(U)\leq-2$.
Then $G[U\cup\{v\}]$ is constructed by deleting an edge $uw$ from $G^{\prime}[U]$, adding a vertex $v$ with $h(v)=2$, and adding two edges $vu$ and $vw$; hence,
$\rho_{G,h}(U\cup\{v\})\leq-2+(2\lambda-1)-2(2\lambda)+(2\lambda+1)=-2,$
contradicting Observation 4.6.
Therefore, $G^{\prime\prime}$ is a copy of $K_{k}$, and for some $\varepsilon\in\{0,1\}$, h(x)=k1h(x)=k-1 for $k-1-\varepsilon$ vertices $x\in U$, and h(y)=k2h(y)=k-2 for $\varepsilon$ vertices $y\in U$. Then, writing $U^{\prime}=U\cup\{v\}$,
$\rho_{G,h}(U^{\prime})=k+(2\lambda-1)-2(2\lambda)+2\lambda-\varepsilon(\lambda+2)=k-1-\varepsilon(\lambda+2).$